OpenStudy (anonymous):

A stationery shop has a sale on address books and calendars. Two of the address books and one calendar cost a total of $10. One of the address books and three calendars cost a total of $15. How much does one calendar cost?

OpenStudy (anonymous):

What is the answer and how you got it?

OpenStudy (anonymous):

Let Address Book = x, Calendars = y 2x + y = 10 ----equation (1) x + 3y = 15 ----equation (2) using eq. (2) x = 15 - 3y ---equation (3) Plugging the value of x in equation (1) 2(15-3y) + y = 10 30 - 6y + y = 10 30 - 5y = 10 -5y = 10 - 30 -5y = -20 y = 4 So the Price of the Calendar is 4 $ each. and using y = 4 in equation 3 x = 15 - 3(4) x = 15 - 12 x = 3 So the price of the Address book is 3$ each. Hope this helps :)

OpenStudy (anonymous):

thank you random awesome person

OpenStudy (anonymous):

Gamer what you play Ghosts :D

OpenStudy (anonymous):

everything lol